OpenStudy (anonymous):

can someone please help me y=x^2-36 y=x-16

Vocaloid (vocaloid):

im guessing you want to solve the system? start by setting the two equations equal to each other x^2 - 36 = x - 16 know where to go from here?

OpenStudy (anonymous):

not really

Vocaloid (vocaloid):

ok, now re-arrange the variables to get it into quadratic form x^2 - x - 20 = 0

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

then factor

OpenStudy (anonymous):

so would that be x^2-36x-20=x

OpenStudy (anonymous):

or is it x^2-36x-20=0

Vocaloid (vocaloid):

nope, remember, to factor, we want to find two numbers that add to -1 and multiply to get -20

Vocaloid (vocaloid):

x^2 - x - 20 = 0 (x-5)(x+4) = 0 so x = 5 and x = -4 now find the matching y values

Vocaloid (vocaloid):

still there?